# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

### v1

@@ Abstract
This paper presents a unified modeling and optimization framework to enhance the kinematic performance of multi-magnet embedded soft continuum robots (MeSCRs). To this end, we establish a differentiable system formulation based on an extended pseudo-rigid-body model. This formulation enables analysis of the equilibrium well-posedness and the geometry of the induced configuration under magnetic actuation. In particular, we show that the maximum controllable degrees of freedom of a MeSCR equal twice the number of embedded magnets. We subsequently develop a structural optimization framework based on differential geometry that links classical kinematic measures (e.g., manipulability and dexterity) to the configuration of embedded magnets. The resulting optimization condition reveals that improving local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion. Closed-form solutions for optimal magnet configurations are derived under representative conditions, and a gradient-based numerical method is proposed for general design scenarios. Simulation studies validate the effectiveness of the proposed framework.
Index Terms:
Magnetic soft continuum robots, pseudo-rigid-body modeling, kinematic performances, structure optimization.
@@ I Introduction
Magnetically actuated soft continuum robots have demonstrated significant potential in medical procedures such as vascular intervention and gastrointestinal examination[1]. Early designs typically embedded a single magnet at the distal end to enable basic magnetic manipulation[2, 3, 4]. However, the limited torque generated by such configurations restricts the controllable degrees of freedom (DoF) less than two[5], significantly constraining spatial maneuverability. Recent advances in magnetic materials and microfabrication have enabled the integration of multiple micro-magnets along the continuum body[6, 7], giving rise to multi-magnet embedded soft continuum robots (MeSCRs). These systems provide multiple distributed sources of actuation torque, substantially enhancing kinematic performance and expanding the design space. Various strategies have been proposed to exploit these capabilities, such as using movable magnets to enlarge reachable workspace[8], designing magnet layouts for multimodal deformation[9, 10], or inducing axial twisting via circumferential arrangements[11]. These efforts can be seen as structural design choices, analogous to tuning joint orientations or link lengths in rigid manipulators to improve kinematic capabilities.
Despite this progress, the absence of a unified modeling framework for MeSCRs remains a major barrier. Unlike rigid-link manipulators, which benefit from well-established analytic tools for performance analysis and optimization, soft continuum robots lack a theoretical foundation that systematically links structure to kinematic function[12]. As a result, many mature strategies developed in rigid manipulators cannot be directly applied to MeSCRs, motivating the need for a new modeling paradigm that bridges this gap.
Current MeSCR modeling approaches typically fall into two categories: Cosserat rod models and pseudo-rigid-body (PRB) models. The Cosserat formulation provides high-fidelity representations and has been integrated with magnetic dipole actuation models for motion control[13, 14]. However, it requires solving boundary value problems derived from variational principles, making it analytically intractable and computationally intensive for iterative design. By contrast, the PRB model discretizes flexible structures into rigid links with elastic joints[15, 16], yielding a Lagrangian formulation analogous to rigid manipulators. This enables unified expressions for forward kinematics and simplifies equilibrium computation via unconstrained finite-dimensional optimization[17]. Early applications in single-magnet MeSCRs have demonstrated effective Jacobian-based control strategies[18, 19], and rank-based analysis has shown that such systems can control at most two DoF[5]. Recent work has extended the PRB formulation to multi-magnet settings[20, 10], preliminarily validating its capacity to simulate complex deformations.
Nevertheless, the PRB model also suffers from several inherent limitations. The choice of joint discretization entails a trade-off between modeling fidelity and computational cost. In particular, due to the lack of theoretical analysis on well-posedness, existing methods typically use numerical methods to analyze the multiplicity and jump phenomena of solutions[21]. Moreover, to our best knowledge, complete closed-form expressions for Jacobian (simplified expression can be found in [20, 19]) and Hessian matrices under multi-magnet configurations have not been established.
To address the limitations outlined above, the first part of this paper investigates an extended-pseudo-rigid-body (EPRB) model[17, 19], which introduces serial spherical joints to enable unified modeling of spatial bending and torsional responses. The EPRB model possesses a generalized expressive capacity: when the elastic modulus approaches zero, it naturally degenerates into a classical rigid-link manipulator model. Conversely, as the number of embedded magnets tends to infinity, it approximates the behavior of a magnetic continuum[22]. Based on this formulation, we conduct a well-posedness analysis of the static equilibrium equations, encompassing uniqueness, boundedness, and convergence properties. Furthermore, we derive complete analytical expressions for the Jacobian and Hessian matrices under multi-magnet configurations, thereby providing theoretical support for both kinematic analysis and controller design.
Building upon the proposed EPRB model, the second part of this paper further demonstrates its ability to facilitate geometric analysis and structural optimization of soft continuum robots. We begin by rigorously proving a key structural result: the maximum controllable degrees of freedom in a magnetically actuated soft continuum robot is exactly twice the number of embedded magnets. This finding corroborates prior empirical observations reported in many existing literature[5, 20]. Additionally, we show that when magnets are arranged axially, the EPRB model naturally reduces to the classical PRB formulation, and a closed-form solution to the equilibrium configuration can be obtained under small-angle premises. As structural optimization often aims to enhance a robot’s kinematic performance, we draw inspiration from rigid manipulator design, where Jacobian-based kinematic measures (e.g., manipulability and dexterity) are widely employed to guide structural synthesis and task planning[23, 24]. We develop a differential geometry-based optimization framework around kinematic measures, which connects kinematic performances to the configuration of embedded magnets in MeSCR. Unlike rigid-link manipulators, whose structural and performance optimization can be directly formulated in joint space, soft continuum robots exhibit implicitly coupled configurations governed by elastomagnetic equilibrium. As such, optimization must be performed over the actuation or design space, with performance indices reparametrized through the configuration manifold. To resolve this, we reinterpret the equilibrium set as a smooth submanifold and apply tools from geometric measure theory to reparametrize global performance objectives over the actuation space. We derive closed-form optimal magnet configuration in representative cases and develop an efficient gradient-based numerical method for the general scenario.
Throughout the technical description, vectors and matrices are defined using lowercase and uppercase bold symbols, respectively, whereas scalars are not bolded.
II Modeling
This section introduces the extended pseudo-rigid-body model to describe the MeSCR. Complete closed-form expressions for the manipulator Jacobian and Hessian matrices are derived. We further investigate the well-posedness of the equilibrium solution. These results provide the necessary mathematical foundation for subsequent kinematic analysis and performance optimization.
II-A Forward Kinematics and Manipulator Jacobian
Figure 1: Illustration of the basic concept of the EPRB model on describing the MeSCR. The continuous body is discretized by a series of rotational rigid rods with joints and frames attached. The bending and twisting deformations of the MeSCR are described by three joint variables $\theta_{\mathbf{t}}^{i}$, $\theta_{\mathbf{u}}^{i}$, and $\theta_{\mathbf{v}}^{i}$.
@@ II-A Forward Kinematics and Manipulator Jacobian
To meet the demands of entering narrow spaces and achieving complex deformations, MeSCRs are typically designed with slender structures. Certain well-founded assumptions contribute to simplifying the model[25], such as the rod remaining free of shear deformation and its cross-section always being orthogonal to the centerline. Additionally, some classical PRB models[15] neglect both twisting and stretching of the continuum. In contrast, we assume only that the continuum is inextensible, with twisting characterized by extended variables [26]. Within the EPRB model framework, a continuum robot of length $L$ is described by a series of rigid rods. As shown in Fig. 1, $N$ spherical joints are distributed along the centerline of the MeSCR at points of interest, connected by $N$ rigid rods with lengths $l_{i},i=0,\dots,N-1$. We use the Voronoi region of the joint to scale the potential energies, which is defined as $\ell_{i}=(l_{i-1}+l_{i})/2,i=0,\dots,N-1$. An orthogonal local frame $\{\bar{\mathbf{t}}_{i},\bar{\mathbf{u}}_{i},\bar{\mathbf{v}}_{i}=\bar{\mathbf{t}}_{i}\times\bar{\mathbf{u}}_{i}\}$ is attached to each joint, with overbars indicating that the MeSCR is in a straightened reference state, and the tangent vector $\bar{\mathbf{t}}_{i}$ remains aligned with the direction of the successive rod. The local frame is used to determine the rotation axis for each joint, expressed as $\boldsymbol{\theta}_{i}=\theta_{\mathbf{t}}^{i}\bar{\mathbf{t}}+\theta_{\mathbf{u}}^{i}\bar{\mathbf{u}}+\theta_{\mathbf{v}}^{i}\bar{\mathbf{v}}\in\mathbb{R}^{3}$, where $\theta_{\mathbf{t}}^{i},\theta_{\mathbf{v}}^{i},\theta_{\mathbf{u}}^{i}\in\mathbb{R}$ are three independent joint variables. The associated rotation matrix is defined as $\mathbf{R}_{i}(\boldsymbol{\theta}_{i})=\mathbf{e}^{[\boldsymbol{\theta}_{i}]_{\times}}\in\mathrm{SO}(3)$, and we define a useful shorthand notation $(\mathbf{R}_{i}^{j})$ to represent the product of rotation matrices that $\mathbf{R}_{i}^{j}=\mathbf{R}_{i}\mathbf{R}_{i+1}\cdots\mathbf{R}_{j}$. With the notation, local frame axes in a deformed state (②), such as the tangent vector, can be easily computed by $\bar{\mathbf{t}}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}$ from the reference state (①). The resulting elastic strain on the $i$th joint will be used to compute the elastic potential energy later.
The configuration space $(\boldsymbol{\theta})$ of the MeSCR can be written as
$\boldsymbol{\theta}=\left[\boldsymbol{\theta}_{0},\dots,\boldsymbol{\theta}_{N-1}\right]^{\top}\in\mathbb{T}\subset\mathbb{R}^{3N_{m}}.$ (1)
@@
Let a fixed base frame $\{s\}$ attach to joint 00, frame $\{b\}$ attach to the distal end, and ${}_{s}^{b}\bar{\mathbf{H}}\in\mathrm{SE}(3)$ be the associated distal pose when the MeSCR is in its zero position. The screw axes in the fixed base frame are of the form:
$\boldsymbol{\xi}_{i}=\begin{bmatrix}\boldsymbol{\theta}_{i}\\ -\boldsymbol{\theta}_{i}\times\bar{\mathbf{p}}_{i}\end{bmatrix}$
@@
where $\bar{\mathbf{p}}_{i}$ is the reference position of the $i$th joint associated with $l_{i}$. Then, the forward kinematics of the MeSCR can be represented in the following product of exponentials:
${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})=\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\mathbf{e}^{\boldsymbol{\xi}_{1}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}{}_{s}^{b}\bar{\mathbf{H}}(\mathbf{0})\in\mathrm{SE}(3).$
@@
We next present the analytical expression of the space Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta})\in\mathbb{R}^{6\times 3N}$, defined in fixed frame coordinates. The space Jacobian relates small changes in joint space to distal pose as $\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\delta\boldsymbol{\theta}$, where $\mathbf{h}=\mathrm{Log}\left({}_{s}^{b}\mathbf{H}\right)\in\mathbb{R}^{6}$ represents the distal pose in Cartesian coordinates. To facilitate calculation, the Jacobian matrix is partitioned into blocks, with each block corresponding to a joint space group as
@@
with $\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}=\begin{bmatrix}\mathbf{I}_{3},\ -[\bar{\mathbf{p}}_{i}]_{\times}\end{bmatrix}^{\top}$, and $\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}$ is the left Jacobian of the $\mathrm{SE}(3)$ manifold, as discussed in [27]. To conclude, we note the following relationship between the space Jacobian and the body Jacobian $(\mathbf{J}_{\boldsymbol{\theta}}^{b}(\boldsymbol{\theta}))$ in case of need:
@@
Also for brevity, $\mathbf{J}_{\boldsymbol{\theta}}$ will be referred to $\mathbf{J}_{\boldsymbol{\theta}}^{s}$ unless specified.
@@ II-B Equations of Motion
Figure 2: Illustration of a quasi-static state of the MeSCR under an external magnetic field $\mathbf{b}(\mathbf{p}_{k})$. Multiple embedded micromagnets at corresponding joints generate magnetic torques $\mathbf{m}_{k}\times\mathbf{b}_{k}$ that induce deformation in the MeSCR. When the magnetic torque and elastic moments reach the equilibrium, the MeSCR attains the quasi-static state.
The MeSCR’s equations of motion are derived from the Lagrangian mechanics, which are formulated based on the MeSCR’s potential energy. As shown in Fig. 2, the MeSCR is allowed to bend, twist, and magnetic actuatable. Thus, the total potential energy features elastic strains and magnetic moments, which are decomposed into the elastic $(\mathcal{E}_{e})$ and magnetic $(\mathcal{E}_{m})$ potential energies as
$\mathcal{E}=\mathcal{E}_{e}+\mathcal{E}_{m}.$ (4)
The respective energies are formulated as follows.
II-B1 Elastic Potential Energy

### v2

# Enhancing Kinematic Performances of Soft Continuum Robots for Magnetic Actuation

@@ Abstract
Soft continuum robots achieve complex deformation through elastic equilibrium, making their reachable motions governed jointly by structural design and actuation-induced mechanics. This work develops a general formulation that integrates equilibrium computation with kinematic performances by evaluating Riemannian Jacobian spectra on the equilibrium manifold shaped by internal/external loading. The resulting framework yields a global performance functional that directly links structural parameters, actuation inputs, and the induced configuration space geometry. We apply this general framework to magnetic actuation. Analytical characterization is obtained under weak uniform fields, revealing optimal placement and orientation of the embedded magnet with invariant scale properties. To address nonlinear deformation and spatially varying fields, a two-level optimization algorithm is developed that alternates between energy based equilibrium search and gradient based structural updates. Simulations and physical experiments across uniform field, dipole field, and multi-magnet configurations demonstrate consistent structural tendencies: aligned moments favor distal or mid-distal solutions through constructive torque amplification, whereas opposing moments compress optimal designs toward proximal regions due to intrinsic cancellation zones.
@@ I Introduction
Soft continuum robots have gained growing attention for tasks involving compliant interaction, dexterous access, and safe manipulation in complex or confined environments. Their ability to realize smooth, multi-segment deformation without rigid joints supports applications in minimally invasive navigation, inspection, and human-centered tasks. Recent advancements have significantly improved perception, control, and structural adaptability in soft continuum systems. Examples include proprioceptive designs capable of simultaneous shape estimation and contact detection [1], learning based deformation estimation under multi-point interaction [2], optimal control strategies for rotation-coupled bending [3], and variable curvature mechanisms that enhance configurability [4]. These developments collectively broaden the functional space of soft continuum manipulation and motivate new models that can capture their nonlinear, compliant behavior.
Magnetic actuation has emerged as a particularly promising approach for soft continuum robots. Compared with tendon-driven or pneumatic mechanisms, magnetic actuation allows wireless control, allows miniaturization to submillimeter scales, and reduces onboard complexity. Recent studies demonstrate substantial gains in workspace, dexterity, and navigation capability across diverse magnetic architectures. Robots with distributed magnetization patterns have shown enhanced deformation diversity [5], movable opposite magnets have enabled workspace expansion [6], and rotating field-assisted designs have achieved buckling free insertion in tortuous lumens [7]. High dexterity helical magnetic robots further extend steerability in endovascular navigation [8]. In parallel, modeling frameworks for magnetic and elastic coupling have evolved rapidly. Analytical and semi-analytical models have been introduced for large deformation magnetoelastic behavior, including kinetostatic formulations capable of capturing multistage bending under complex dipole fields [9], modular magnetization profile evaluation [10], and optimized structural patterns for enhanced deformability. Additional works explore interactions between magneto elasticity, external loading, and nonlinear geometries, providing foundations for real-time control and structural tuning. Together, these contributions highlight the importance of accurate equilibrium modeling and structural design for improving kinematic performance.
However, several core challenges remain unresolved. Magnetic soft continuum robots do not operate on a fixed kinematic map. Their attainable configurations are determined by the equilibrium manifold produced by the interaction between elastic restoring forces and magnetic loading. This manifold is highly nonlinear and sensitive to magnet placement, field variation, torque cancellation, and bifurcation events. Existing design methods often rely on task-specific heuristics [6], empirical tuning of magnet distributions [5], data-driven control [11, 12], or simplified kinematic approximations that overlook the intrinsic coupling between mechanics and actuation. Learning-based or adaptive controllers [13, 1] enhance robustness, yet they offer limited insight into the structural sources underlying the equilibrium geometry. Moreover, commonly used performance indices such as manipulability and dexterity, being defined in Euclidean coordinates, may only partially capture the curvature of the equilibrium manifold under magnetic loading. As a result, structural optimization and performance evaluation are often decoupled, limiting the ability to systematically ensure actuation feasibility, stability, and kinematic quality.
These gaps motivate the need for a unified formulation that couples equilibrium with kinematic performance. Since the robot’s motion is realized only through equilibrium configurations, performance should be evaluated on the manifold shaped by internal and external loading. It further requires a structural optimization scheme capable of handling nonlinear magnetoelastic interactions, diverse field geometries, and multi-magnet architectures recently explored in continuum and catheter-like robots [14, 6, 7].
In this study, we develop a unified framework and demonstrate its theoretical and practical utility across homogeneous field and dipole field actuation. We leverage Riemannian Jacobian spectra to characterize manipulability and dexterity on the equilibrium manifold and integrate these indices with an energy-based equilibrium computation. This formulation enables physically consistent performance evaluation and reveals structural trends that persist across actuation regimes, including robots with modular magnetization segments [10], multi-magnet bodies [5], and clinically oriented continuum tools [15, 16]. The main contributions of this work are summarized as follows:
  1. 1.
    We establish a unified structural optimization framework that couples magnetic energy-based equilibrium analysis with Riemannian Jacobian spectral evaluation on the equilibrium manifold. By formulating the equilibrium set as an actuation-dependent submanifold and computing kinematic performance on its pullback metric.
  2. 2.
    We derive analytical results for soft continuum robots actuated by weak uniform magnetic fields, including closed-form optimality conditions for magnet placement and orientation. These theoretical characterizations reveal scale-invariant structural principles that are further validated through numerical optimization.
  3. 3.
    We extend the proposed framework to general dipole-generated magnetic fields and apply the full two-level optimization scheme to multi-magnet configurations. Comprehensive numerical studies and corresponding physical experiments demonstrate consistent trends in equilibrium behavior and kinematic performance across diverse actuation paths.
The remainder of this paper is organized as follows. Sec. II presents the modeling framework, including the pseudo-rigid-body representation, magnetic torque formulation, and energy-based equilibrium analysis. Sec. III defines the kinematic performance metrics and formulates the structural optimization problem. Sec. IV introduces the analytical and numerical optimization methods. Sec. V provides case studies and numerical validations under uniform and general magnetic fields. Finally, Sec. VI discusses the implications of the proposed framework and concludes the paper. Throughout the technical description, vectors and matrices are defined using lowercase and uppercase bold symbols, respectively, whereas scalars are not bolded.
II Kinematics of Soft Continuum Robots
II-A Kinematic Mapping
Figure 1: Illustration of the basic concept of the PRB model on describing the SCR. The soft body is discretized into rigid rods connected by joints with local frames. The configuration is represented by the joint variables $\theta_{\mathbf{t}}^{i}$, $\theta_{\mathbf{u}}^{i}$, and $\theta_{\mathbf{v}}^{i}$. The system reaches an equilibrium state (②) when the torques of the actuators (①) are balanced by the elastic restoring moments, and the joint deformation is stored as elastic strain.
We consider SCRs with slender structures, whose length is significantly greater than their radius. Certain well-founded assumptions contribute to simplifying the model [17], such as the rod remaining free of shear deformation and its cross-section always being orthogonal to the centerline. Additionally, some classical PRB models[18] neglect both twisting and stretching of the continuum. In contrast, we assume only that the continuum is inextensible, with twisting characterized by extended variables [19].
@@ II-A Kinematic Mapping
Within the PRB model framework, the SCR with length $L$ is described by a series of rigid rods. As shown in Fig. 1, $N$ spherical joints are distributed along the centerline of the SCR at points of interest, connected by $N$ rigid rods with lengths $l_{i},i=0,\dots,N-1$. We use the Voronoi region of the joint to scale the potential energies, which is defined as $\ell_{i}=(l_{i-1}+l_{i})/2,i=0,\dots,N-1$. An orthogonal local frame $\{\bar{\mathbf{t}}_{i},\bar{\mathbf{u}}_{i},\bar{\mathbf{v}}_{i}=\bar{\mathbf{t}}_{i}\times\bar{\mathbf{u}}_{i}\}$ is attached to each joint, with overbars indicating that the SCR is in a straightenedl reference state, and the tangent vector $\bar{\mathbf{t}}_{i}$ remains aligned with the direction of the successive rod. The local frame is used to determine the rotation axis for each joint, expressed as $\boldsymbol{\theta}_{i}=\theta_{\mathbf{t}}^{i}\bar{\mathbf{t}}+\theta_{\mathbf{u}}^{i}\bar{\mathbf{u}}+\theta_{\mathbf{v}}^{i}\bar{\mathbf{v}}\in\mathbb{R}^{3}$, where $\theta_{\mathbf{t}}^{i},\theta_{\mathbf{v}}^{i},\theta_{\mathbf{u}}^{i}\in\mathbb{R}$ are three independent joint variables. The associated rotation matrix is defined as $\mathbf{R}_{i}(\boldsymbol{\theta}_{i})=\mathbf{e}^{[\boldsymbol{\theta}_{i}]_{\times}}\in\mathrm{SO}(3)$, and we define a useful shorthand notation $(\mathbf{R}_{i}^{j})$ to represent the product of rotation matrices that $\mathbf{R}_{i}^{j}=\mathbf{R}_{i}\mathbf{R}_{i+1}\cdots\mathbf{R}_{j}$. With the notation, local frame axes in a deformed state (②), such as the tangent vector, can be easily computed by $\mathbf{t}_{i}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}$ from the reference state (①). The resulting elastic strain on the $i$th joint will be used to compute the elastic potential energy later.
The configuration space $(\boldsymbol{\theta})$ of the SCR can be written as
$\boldsymbol{\theta}=\left[\boldsymbol{\theta}_{0},\dots,\boldsymbol{\theta}_{N-1}\right]^{\top}\in\mathbb{T}\subset\mathbb{R}^{3N}.$ (1)
@@
Let a fixed base frame $\{s\}$ attach to joint 0, frame $\{b\}$ attach to the distal end, and $_{b}^{s}\bar{\mathbf{H}}\in\mathrm{SE}(3)$ be the associated distal pose when the SCR is in its zero position. The screw axes in the fixed base frame are of the form
$\boldsymbol{\xi}_{i}=\begin{bmatrix}\boldsymbol{\theta}_{i}\\ -\boldsymbol{\theta}_{i}\times\bar{\mathbf{p}}_{i}\end{bmatrix}\in\mathbb{R}^{6}$
@@
where $\bar{\mathbf{p}}_{i}$ is the reference position of the $i$th joint associated with $l_{i}$. Then, the forward kinematics of the SCR can be represented in the following product of exponentials:
$_{b}^{s}\mathbf{H}(\boldsymbol{\theta})=\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\mathbf{e}^{\boldsymbol{\xi}_{1}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}{}_{b}^{s}\bar{\mathbf{H}}(\mathbf{0})\in\mathrm{SE}(3).$ (2)
@@
We next present the analytical expression of the space Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta})\in\mathbb{R}^{6\times 3N}$, defined in fixed frame coordinates. The space Jacobian relates small changes in joint space to distal pose as $\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\delta\boldsymbol{\theta}$, where $\mathbf{h}=\mathrm{Log}\left({}_{b}^{s}\mathbf{H}\right)\in\mathbb{R}^{6}$ represents the distal pose in Cartesian coordinates. The Jacobian matrix is partitioned into blocks, with each block corresponding to a joint space group:
@@
with $\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}=\begin{bmatrix}\mathbf{I}_{3},\ -[\bar{\mathbf{p}}_{i}]_{\times}\end{bmatrix}^{\top}$, and $\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}$ is the left Jacobian of the $\mathrm{SE}(3)$ manifold as discussed in [20]. To conclude, we note the following relationship between the space Jacobian and the body Jacobian $(\mathbf{J}_{\boldsymbol{\theta}}^{b}(\boldsymbol{\theta}))$ in case of need:
@@
Also for brevity, $\mathbf{J}_{\boldsymbol{\theta}}$ will be referred to $\mathbf{J}_{\boldsymbol{\theta}}^{b}$ unless specified.
@@ II-B Equations of Motion
The equations of motion are derived within the framework of Lagrangian mechanics, formulated in terms of the potential energies stored in the SCR. The SCR is assumed to undergo bending and twisting deformations and to be driven by generalized actuation sources, which can represent either distributed or localized internal/external torques. Accordingly, the total potential energy is composed of elastic strains and actuation-induced moments, decomposed into the elastic $(\mathcal{E}_{e})$ and actuation $(\mathcal{E}_{u})$ energies:
$\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u})=\mathcal{E}_{e}(\boldsymbol{\theta})+\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u}),$ (5)